OpenStudy (ashsmashum):

write an equation in point slope form of the line that passes through the given point and with the given slope m (4,2);m=-5/3

OpenStudy (sunnnystrong):

Reca\[y-y_{1}=m(x-x_{1})\]ll that: Point slope form--> @ashsmashum what do you think?

OpenStudy (ashsmashum):

so 4-5 and 2-3?

OpenStudy (sunnnystrong):

@ashsmashum if (4,2)= (x1,y1) & m= -5/3 than \[y-2=-\frac{ 5 }{ 3 }(x-4)\]

OpenStudy (saov):

Let me help you, x1 and y1 refer to any point (x,y) of the line, because you can choose x1 = 4 , and y1 = 2

OpenStudy (ashsmashum):

so 4-1 and 2-1 or 5-1 and 3-1

OpenStudy (saov):

what do you mean about 4-1 , 2-1 , ..?

OpenStudy (ashsmashum):

I'm sorry if I'm wrong i am very bad at math

OpenStudy (sunnnystrong):

@ashsmashum ... don't worry about it! everyone has to learn this stuff.

OpenStudy (saov):

Don`t worry, you can learning by practice and practice

OpenStudy (jdoe0001):

hint: \(\bf \begin{array}{lllll} &x_1&y_1\\ % (a,b) &({\color{red}{ 4}}\quad ,&{\color{blue}{2}})\quad \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies -\cfrac{5}{3} \\ \quad \\ % point-slope intercept y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (ashsmashum):

but what is x1 and y1?